balance the following nuclear reaction
3 pts Question 29 Balance the following nuclear reaction. 4 327 Np He + 93 2
Answer #1

Step 1

Determine the missing mass number

mass number of parent atom = 4 + 237 = 241

Step 2

Determine missing atomic number

Atomic number of parent atom = 2+93 = 95.

Hence parent atom is Americium ( Am).

So, balance nuclear reaction is

241Am 95\large \rightarrow4 He 2 + 237 Np 93

[Note: Mass number of Np is 237. Not 327. ]